OpenStudy (anonymous):

Solve the system of equations. 3x – 7y = –13 –2x + 3y = 7

OpenStudy (cwrw238):

hint: multiply first equation by 2 and second by 3

OpenStudy (sleepyjess):

3x -7y + 7y = 7y + 13 3x = 7y + 13 \[x = \frac{ 7y+13 }{ 3 }\]

OpenStudy (anonymous):

use substitution.

OpenStudy (sleepyjess):

Then substitute x into the 2nd equation

OpenStudy (sleepyjess):

\[-2(\frac{ 7y+13 }{ 3 })+3y=7\]